OpenStudy (anonymous):

Find the radius, if the angle is 90degrees and the arc is 104

OpenStudy (anonymous):

i'm guessing this is a circle with an arc length of 104?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

if so then the 104 is a fourth of the circumference circum = 2pi*r

OpenStudy (anonymous):

circum = 4*104

OpenStudy (anonymous):

416?

OpenStudy (anonymous):

ya then just solve for r

OpenStudy (anonymous):

so than divide 416 by

OpenStudy (anonymous):

r = 208/pi

OpenStudy (anonymous):

66.24

OpenStudy (anonymous):

arc=angle * r [104=(\pi/2)*r\] \[r=(2*104)\div(\pi)\] \[r=208/\pi\]